ganeshie8 (ganeshie8):

It is given that \(n\) is not a perfect cube. If \(\prod\limits_{d\mid n} d = n^2\), then find the value of \(\prod\limits_{d\mid n^2}d\) Options : a) n^3 b) n^4 c) n^6 d) n^8 e) n^9

OpenStudy (anonymous):

is this the answer->n^4 ?

ganeshie8 (ganeshie8):

nope

ganeshie8 (ganeshie8):

may i see your work ? it could be some arithmetic error...

ganeshie8 (ganeshie8):

\(\prod\limits_{d\mid n} d\) means : "product of all the positive divisors of n"

ganeshie8 (ganeshie8):

example : n = 10 \(\prod\limits_{d\mid 10} d = 1*2*5*10 =100 \)

OpenStudy (pawanyadav):

I got it

OpenStudy (anonymous):

All numbers that divide n must also divide n^2 and i realized what mistake i made i thought the greatest divisor of n^2 other than it will be n and then for \(\prod\limits_{d\mid n^2}d\) we just have an extra n^2 so we multiply it with\(\prod\limits_{d\mid n} d \) and get n^4 but my new approach-> is a divides n then a^2 will divide n^2 so divisors of n^2 will be all the divisors of n+ the square of all these divisors so \[\prod\limits_{d\mid n^2} d =\prod\limits_{d\mid n} d \times \left( \prod\limits_{d\mid n} d \right)^2\]\[\prod\limits_{d\mid n^2} d=n^2 \times n^4\]\[\prod\limits_{d\mid n^2} d=n^6\]

ganeshie8 (ganeshie8):

that is a very nice try, but the answer is still wrong

ganeshie8 (ganeshie8):

consider an example

ganeshie8 (ganeshie8):

\(n = 2*3\) the divisors of \(n\) are \(\{1,2,3,6\}\) the divisors of \(n^2\) are \(\{1,2,3,6,4,12,9,18,36\}\) notice that there are few extra divisors in \(n^2\). In your proof, you're not counting these divisors of \(n^2\) : \(\{12, 18\}\)

OpenStudy (anonymous):

ok ok i got where i was wrong as again if we take some number say n the prime factorization will be like-\[n=2^a 3^b5^c.....\]total factors will be given by all combination of multiplication of the factors so total multiples will be-> \(No.~Of~Multiples~Of~n=(a+1)(b+1)(c+1)....\) if we square n then we get prime factorization like this-\[n^2=2^{2a}3^{2b}5^{2c}.\] \(total~multiples~of~n^2=(2a+1)(2b+1)(2c+1).....\) which is proves that there will be more factors

ganeshie8 (ganeshie8):

when you say "multiiple", you surely mean "divisors" right ?

OpenStudy (anonymous):

sorry yes it is divisors

OpenStudy (anonymous):

so 3 of the options are eliminated :)

OpenStudy (anonymous):

i think we should use the information that n is not perfect cube

ganeshie8 (ganeshie8):

yes

OpenStudy (pawanyadav):

So we got the answer n^9

ganeshie8 (ganeshie8):

that is the correct answer

OpenStudy (pawanyadav):

BTW answer is not so important

OpenStudy (pawanyadav):

The way is more

OpenStudy (pawanyadav):

If we take n=8 which is a perfect cube What will be the problems....

OpenStudy (pawanyadav):

1*2*4*8 are 8 divisor And their product is also 8 square .

OpenStudy (pawanyadav):

Divisor of 64 are 1*2*4*8*16*32*64 And their product is 8^7

OpenStudy (pawanyadav):

That why we are not given with n=perfect cube Because some divisor remains same that's why we get less number of divisor than 36

OpenStudy (anonymous):

i like this method of assumption, we get the answer real quick :D we make proper cases consider\[n=p_1 \times p_2\]where \(p_1~and~p_2\) are primes its clear that n is not a perfect cube the divisors of n will be->\(1,~p_1,~p_2,~n\) so \(\prod\limits_{d\mid n} d = n^2\) holds good now when we see n^2\[n^2=p_1 \times p_2 \times p_1 \times p_2\]we can easily see all diff divisors of n^2 all diff divisors-\(1,~p_1,~p_2,~n,~p_1^2,p_2^2,~p_1^2p_2,~p_1p_2^2,~n^2\) multiplying them we get->\((p_1p_2)(p_1p_2)^{2}(p_1p_2)^3(n)(n^2)\)=\(n^9\)

ganeshie8 (ganeshie8):

brilliant!

OpenStudy (anonymous):

but this is not general solution because i made it up

ganeshie8 (ganeshie8):

you can make it general

ganeshie8 (ganeshie8):

you just need to prove below lemma first \[\large n^{\tau(n)} = \left(\prod\limits_{d\mid n}d \right)^2\]

ganeshie8 (ganeshie8):

that restricts the number of prime factors to 2 as you're doing in your proof

OpenStudy (pawanyadav):

Am also thinking about this method but I can't find the number of prime factor. How can you say that they are only 2 It luckily coincide with 6 which is having only two primfactors.

OpenStudy (dan815):

http://prntscr.com/97lwd2

ganeshie8 (ganeshie8):

that will do :)

ganeshie8 (ganeshie8):

im sure this identity isn't hard for you to prove \[\large n^{\tau(n)} = \left(\prod\limits_{d\mid n}d \right)^2\]

ganeshie8 (ganeshie8):

that conforms that \(n\) must have exactly two prime factors

ganeshie8 (ganeshie8):

\(\tau(n)\) = number of positive divisors of \(n\)

OpenStudy (dan815):

oh thats neat

ganeshie8 (ganeshie8):

\[\large n^{\tau(n)} = \left(\prod\limits_{d\mid n}d \right)^2\] take srt both sides and get \[\large n^{\tau(n)/2} =\prod\limits_{d\mid n}d \tag{1}\] we're given \(\prod\limits_{d\mid n}d = n^2\tag{2}\) comparing (1) and (2) gives \[\tau(n)=4\]

OpenStudy (dan815):

how to generalize

ganeshie8 (ganeshie8):

\(\tau(n)=4\) means \(n\) is of form \(p^3\) or \(p_1p_2\)

ganeshie8 (ganeshie8):

because the number of divisors of \(n=p_1^{e_1}p_2^{e_2}\cdots p_r^{e_r}\) is given by \((e_1+1)(e_2+1)\cdots(e_r+1)\)

ganeshie8 (ganeshie8):

we can get \(\tau(n)=4\) only in two ways : (3+1) (1+1)(1+1)

OpenStudy (dan815):

ok yup

ganeshie8 (ganeshie8):

and we're given that \(n\) is not a perfect cube, so we can conclude \(n\) is a product of two distinct primes

OpenStudy (dan815):

im okay with that part, i mean how to generalize for all t(n)

ganeshie8 (ganeshie8):

rest is your/elusiveavenger's proof

OpenStudy (dan815):

i can see an induction way to build up a multiple of cases, or try and work out t(n)=k d|k of form (e1+1)(e2+1)....(en+1)

OpenStudy (dan815):

this would give us how many perfect squares and distinct primes we are dealing with

OpenStudy (dan815):

but its just expressions not closed form, the general case for all t(n) looks like work

ganeshie8 (ganeshie8):

we have showed that the given statement works only when \(n\) is a product of two distinct primes, right ?

OpenStudy (dan815):

ya

ganeshie8 (ganeshie8):

can u explain a bit more... what generalization ... i feel im not seeing something.. .

OpenStudy (dan815):

like this identity we saw to be true for t(n)=4 for once case, and that was because we saw how to move the d/n^2 case

OpenStudy (dan815):

but we still have to prove it really works for alll possible t(n) cases dont we

ganeshie8 (ganeshie8):

Oh you're referring to this identity ? \[\large n^{\tau(n)} = \left(\prod\limits_{d\mid n}d \right)^2\]

ganeshie8 (ganeshie8):

i don't have a proof yet... need to think a bit... but im sure it wont be too hard..

ganeshie8 (ganeshie8):

il open a new question for this proof it looks a bit interesting

OpenStudy (dan815):

okk

OpenStudy (kainui):

\[\frac{\tau(n)}{2} = 2\] n has 4 divisors, in other words \(n=p*q\) or \(n=p^3\) BUT n is not a perfect cube!